OpenStudy (anonymous):

Medal and Fan Are the ratios 25/45 and 15/27 proportional?

OpenStudy (anonymous):

I think they are, but im not 100% sure. sorry

OpenStudy (anonymous):

yes they are

OpenStudy (mrbaconqwerty):

They are

OpenStudy (anonymous):

Can someone explain why please <3

OpenStudy (anonymous):

if you simplify them you will be able to see

OpenStudy (anonymous):

Ok thx guys!

OpenStudy (anonymous):

you are so welcome

OpenStudy (texaschic101):

check them... 25/45 = 15/27 --- cross multiply (25)(27) = (45)(15) now do the math...if both sides come out equal, then they are proportional. If they do not come out equal, then they are not proportional.

OpenStudy (anonymous):

Thx!

OpenStudy (anonymous):

is this all

OpenStudy (anonymous):

you can fan me and then you can tag me with ather Questions

OpenStudy (anonymous):

I have one more I'm going to post it

OpenStudy (cwrw238):

25 / 45 = 5 /9 (dividing top and bottom by 5 15 / 27 = 5/9 ( dividing by 3)

OpenStudy (anonymous):

ok